OpenStudy (anonymous):

WRITE THIS RATIO AS FRACTION IN LOWEST TERM $16.32 TO $21.28 IS ..................(SIMPLIFY YOUR ANSWER. TYPE A FRACTION)

OpenStudy (anonymous):

Might as well just take out the decimal for now and write it as \[\large \sf \frac{1632}{2128}\] Then just simplify as far as you can.

OpenStudy (anonymous):

UMM I STILL LOST

OpenStudy (anonymous):

The easiest way to simplify bigger stuff is to keep going down by 2's whenever both numbers are even. \[\large \sf \frac{1632}{2128} \div \frac{2}{2}\]

OpenStudy (haseeb96):

look as you get 1632/2128 now divide both terms by 2 and  and you get 408/507